# SPrimary Studies

### v1

@@ 3 State of the art
More recently, Andrade et al. [13] investigated the misuse of personal data and what this entails in terms of individuals’ privacy and software product’s quality. The authors analyzed 75 papers to understand how PbD principles had been applied in the SE practices. The results highlighted the lack of specific methodology and tools for translating these principles into practical activities throughout the software development lifecycle. This concern has become even more relevant now that PbD is part of regulations such as GDPR.
More related to our work, Alloghani et al. [23] focused on identifying security and privacy issues in mobile devices and systems alongside existing methods for detecting and preventing such issues. The authors emphasized the importance of utilizing preventive, detective, and responsive methods for ensuring security and privacy in mobile applications. Such methods require the involvement of both mobile vendors and service providers.
Shrivastava et al. [25] analyzed privacy issues concerning permission requests in Android apps. The authors studied 110 research papers, identifying possible severe security implications of the Android permission protocol and further describing several limitations in permission checks.
Ebrahimi et al. [33] analyzed 59 papers related to privacy in mobile apps. They focused on the practice of obtaining explicit consent prior to accessing sensitive private information, in exchange for offering a more personalized user experience. The authors argue that such privacy-invading practices have led to major privacy concerns among app users. The authors also showed that existing literature had focused mainly on detecting data leaks, with little to no attention being given to the user’s perspective.
Negri-Ribalta et al. [34] surveyed the impact of GDPR’s on regulatory requirements pertinent to data protection and how they are handled from an RE perspective. Their study involved 90 papers published between 2016 (i.e., right about when the regulation was voted) and 2022 (four years after GDPR was actually enforced). The work provides a generic view of the GDPR provisions. For the analysis, the authors created a taxonomy of GDPR privacy concerns according to an ad-hoc manual analysis of the regulation. In comparison, our work is driven by a comprehensive and detailed conceptualization of privacy-relevant concerns in GDPR. Mapping existing studies in the literature against this conceptual model provides guidance for future comparisons.
To conclude, in contrast with the aforementioned secondary studies, in this SLR we present a comprehensive overview of the existing research landscape in SE, focusing on privacy in mobile apps. Our SLR further projects the current research onto a reference model of GDPR privacy concepts.
@@ 4.1 Research Questions (RQs) Definition
RQ1. How well does the literature on privacy in mobile apps cover the GDPR privacy metadata types? This RQ explores the research landscape on mobile apps to identify the main topics in terms of GDPR privacy requirements. Drawing on the reference model (presented in Section 2), we answer RQ1 by mapping the topics investigated in existing work on mobile apps to the metadata types in the conceptual model. We further outline the GDPR principles that are investigated in the literature but not represented in the conceptual model.
RQ2. What are the main objectives pursued when investigating GDPR privacy concepts? RQ2 explores the main objectives of primary studies that investigated GDPR privacy in mobile apps. The intuition behind this RQ is to understand the SE research problems related to GDPR that were acknowledged or addressed by existing work. This helps to understand which problems or GDPR privacy metadata remain under-investigated, leading to potentially open directions for future research.
RQ3. What are the different types of research contribution? RQ3 explores the different types of research contributions made by the primary studies, distinguishing between technical and non-technical contributions, and analyzing the app stores and operating systems targeted in the literature. It provides a comprehensive view on the research contributions and further describes, in the case of technical solutions, the enabling technologies and the way such contributions were validated.
RQ4. Are the research artifacts accompanying studies on GDPR privacy concepts publicly available? This RQ sheds light on the publicly available artifacts that can be potentially leveraged in future research. RQ4 further reports on the licenses under which these artifacts have been shared.
@@ 4.2 Phase I: Paper Extraction
Ensuring the retrieval of a relatively representative sample of the existing literature is essential for conducting an SLR. This phase involves three steps: (1) selecting the right databases or search engines to query, (2) appropriately refining the query keywords, and (3) expanding our results by retrieving additional papers from indexing systems. Next, we elaborate on each step.
@@ 4.2.1 Databases Selection
For the search of papers in the different databases, we adopted common practices for and recommendations from the SLR literature [37, 36, 38]. Specifically, we retrieved primary studies from the following databases: ACM Digital Library [^1] https://dl.acm.org/, IEEE Xplore [^2] https://ieeexplore.ieee.org/, ScienceDirect [^3] https://www.sciencedirect.com/, Scopus [^4] https://www.scopus.com/, and SpringerLink [^5] https://link.springer.com/. These sources are typically selected in SLRs and systematic mapping studies [39, 40]. Additionally, we searched for papers in the following search engines or indexing systems: DBLP [^6] https://dblp.org/, Google Scholar [^7] https://scholar.google.com/, and Web of Science [^8] https://www.webofscience.com/wos/. Our motivation to include these additional sources is driven by two reasons. First, we enlarged our search process to ensure a complete coverage of the SE and RE literature, assuming that these sources likely contain other journals and conference proceedings that might be missing from the regular databases. Second, our paper search using regular sources might miss papers due to being, for instance, not yet in their final publisher’s form. We therefore scoped our search against these indexing systems by restricting it to only years 2022 and 2023.
4.2.2 Search Queries Definition
@@
Initially, we tested the following query, which combines several keywords: (“mobile app*” AND “data protection” AND “data privacy”). We then created variants where we used single terms as keywords, e.g., (“mobile” AND “app*”) instead of (“mobile app*”). We queried all databases with the same pre-defined set of keywords except for ScienceDirect, since it does not allow defining queries with special characters such as wildcards(*). In this case, we used the following combination: (“mobile app” OR “mobile apps” OR “mobile application” OR “mobile applications”).
Following this, we iteratively refined our queries. To ensure that the retrieved papers were within the scope of our study, our final queries comprised the following keywords: “mobile app”, “privacy”, and “gdpr”. The final queries are listed in Table 1.
Table 1: Final Search Queries and Respective Number of Papers Retrieved per Database.
[t] Query Database N “mobile apps” AND “privacy” AND “gdpr” ACM Digital Library 181 “mobile apps” AND “privacy” AND “gdpr” IEEE Xplore 70 “mobile apps” AND “privacy” AND “gdpr” ScienceDirect 129 “mobile apps” AND “privacy” AND “gdpr” Scopus 26 “mobile apps” AND “privacy” AND “gdpr” SpringerLink 78 “mobile apps” AND “privacy” AND “gdpr” Web of Science 17 “mobile apps” AND “privacy” AND “gdpr” DBLP 15 “mobile apps” AND “privacy” AND “gdpr” Google Scholar 100 Total All 616
    N: Number of retrieved papers.
@@ 4.2.3 Papers Retrieval
Querying the regular databases, as shown in Table 1, resulted in a collection of 484 papers, distributed across the different databases as follows: 181 from ACM Digital Library, 70 from IEEE Xplore, 129 from ScienceDirect, 26 from Scopus, and 78 from SpringerLink.
As noted earlier, we also queried indexing systems for additional potentially missing papers. Consequently, we retrieved 132 additional papers, including 100 papers from Google Scholar, 17 papers from the Web of Science, and 15 papers from DBLP. As for Google Scholar, we considered only the first 10 pages of results, assuming that papers appearing after that will not likely be relevant to the objectives of our study.
Our paper extraction phase resulted in a total of 616 papers, which were then passed on to the next phase for further analysis.
@@ 4.3.1 Paper filtering
We applied duplicate removal on the entire collection of the 616 papers. To identify duplicates, we defined a case-insensitive Excel macro that compares the titles of the papers and determines possible duplicates. The output of this macro indicated which paper metadata details (particularly the title) were actually duplicates. This output was then manually vetted by the first author of this paper. As a result, we filtered out 204 papers which were deemed genuine duplicates, thereby reducing our collection to 412 papers.
@@
    Paper length: We defined a minimum paper size of six pages (including references). The rationale behind this threshold was to discard short papers such as position/vision/demo/new-idea/early-research-achievements papers with only preliminary or partial results, as commonly done in relevant literature [39, 40].
@@
    Publication venues: Following the guidelines of Kitchenham and Brereton [27], we filtered out papers that were not relevant for our study. Specifically, we excluded papers published in journals or conferences clearly unrelated to SE and security and privacy since privacy has long been studied from the software security perspective. In this way, we removed papers focusing on the social aspects of privacy/data protection or investigating specialized topics. In case of uncertainty (e.g., when a paper title appeared relevant) we retained the venue and related paper(s) for further analysis (see below). Examples of excluded venues are the International Journal of Medical Informatics, Blockchain: Research and Applications, and the EAI International Conference on Smart Objects and Technologies for Social Good.
After applying the paper filtering, our collection was reduced to 168 papers which were further processed in the successive step.
Table 2: Journals and Conference Venues with Number of Papers Deemed Relevant in this SLR
[t] ID Venue N J1 Wireless Networks 1 J2 Frontiers of Computer Science 1 J3 Software and Systems Modeling 1 J4 Computing 1 J5 Empirical Software Engineering 2 J6 AI & Society 1 J7 International Journal of Information Security 1 J8 Personal and Ubiquitous Computing 1 J9 Internet Interventions 1 J10 Computers & Security 4 J11 Journal of Systems and Software 1 J12 Information and Software Technology 1 J13 Computer Law & Security Review 4 J14 ACM Transactions on Computer-Human Interaction 7 J15 ACM Transactions on Software Engineering Methodology 1 J16 IEEE Transactions on Software Engineering 2 V1 IEEE International Conference on Software Analysis, Evolution and Reengineering 1 V2 The Web Conference 1 V3 IEEE/ACM International Conference on Automated Software Engineering 5 V4 Workshop on Privacy in the Electronic Society 1 V5 ACM Conference on Security and Privacy in Wireless and Mobile Networks 1 V6 CHI Conference on Human Factors in Computing Systems 1 V7 ACM/IEEE International Conference on Software Engineering 6 V8 ACM Web Conference 1 V9 Annual Computer Security Applications Conference 1 V10 International Conference on Availability, Reliability and Security 1 V11 ACM SIGSAC Conference on Computer and Communications Security 2 V12 Annual IEEE/IFIP International Conference on Dependable Systems and Network 1 V13 USENIX Security Symposium 2 V14 IEEE International Requirements Engineering Conference 3 V15 IEEE European Symposium on Securityand Privacy Workshops 2 V16 IEEE Annual Computer Software and Applications Conference 1 Total 60
    N: Number of papers deemed relevant in this SLR.
4.3.2 Relevance Assessment
In this step, we analyzed the remaining papers and identified those relevant for our SLR. The first three authors collectively participated in this process. Driven by the motivation of our SLR, the relevance of the papers was assessed according to their main focus being privacy. Relevant papers should clearly discuss GDPR-relevant privacy concepts and how these were handled in mobile apps. For instance, a paper that discusses users practices in mobile apps, such as subscription cancellation and asking for refunds, or users providing inaccurate information were considered as not relevant to our SLR. Similarly, papers discussing dark patterns that attempt to manipulate users’ behaviors are not relevant to our SLR. While these papers are potentially relevant to the overall privacy concept, they are not particularly informative about how GDPR privacy concepts on handling and processing personal data were addressed in mobile apps. This step spanned three iterations, described below.
@@ 4.3.2 Relevance Assessment
This step allowed us to perform a further selection of papers as well as of the relevant journals and venues (which were reduced to 32, equally split—see Table 2). The output of this step is a final set of 60 papers (corresponding to 35.7% of the 168 analyzed papers) on which we conducted our data extraction phase.

### v2

@@ 3.3 Privacy-by-Design (PbD) in SE
Finally, Morales-Trujillo et al. [8] analyzed privacy by design (PbD) in software systems. Their work aimed to identify relevant literature collecting PbD goals (e.g., data protection mechanisms) in software development and determining the extent to which PbD was present in the current software development practices. The authors analyzed 49 papers and concluded that the primary studies mainly focus on data minimization. The authors further suggested that PbD was still an underdeveloped concept in the SE literature and acknowledged the necessity for a framework to support the development of privacy-aware systems.
3.4 Motivation for this work
Table 1 summarizes and compares the aforementioned secondary studies against ours. We note that only two studies have explicitly considered the GDPR, including Negri-Ribalta et al.’s work [38] on general privacy concerns under the GDPR and Saltarella et al.’s survey concerning PbD [41]. The majority of the remaining studies have no explicit legal ground in their work. In contrast, in this SLR we present a comprehensive overview of the existing research landscape in SE, focusing on privacy in mobile apps. Our SLR further projects the current research onto a reference model of GDPR privacy concepts.
Table 1: Analysis of the State of the Art
Grounding Legal Basis Studies Privacy-related Concepts Focus
Sangaroonsilp & Dam [35] N/A 187 13 privacy threats Privacy trends in the SE literature
Showail et al. [36] PDPL 66 5 privacy concerns Saudi Arabia’s Personal Data Protection Law (PDPL) compliance in mobile apps
Ebrahimi et al. [37] N/A 59 N/A Consent and access to personal information in mobile apps
Alloghani et al. [25] N/A 330 N/A Security and privacy in mobile devices and systems
Shrivastava et al. [11] N/A 110 1 (Consent) Security and Privacy of Android Permission System
Negri-Ribalta et al. [38] The GDPR 90 30 Requirements related to the GDPR in the literature
Fernández-Alemán et al. [2] N/A 49 5 high-level concerns Security and privacy in electronic healthcare records
Iwaya et al. [24] NIST-8062 365 20 controls from the standard Security and privacy in mHealth and uHealth systems
Martínez-Pérez et al. [7] N/A 169 7 high-level concerns Security and privacy in mHealth apps
de Chaves et al. [40] N/A 68 7 PbD principles Level of understanding of PbD in the SE literature
Andrade et al. [12] N/A 75 N/A Privacy by design
Saltarella et al. [41] The GDPR 91 7 PbD principles Best practices of design and development for the PbD and privacy-by-default paradigms
Semantha et al. [10] N/A 23 7 PbD principles Privacy by design in healthcare
Morales-Trujillo et al. [8] ISO/IEC 29100 49 11 ISO/IEC 29100 principles Privacy by design
This work The GDPR 60 56 GDPR-related Privacy in Mobile Apps
  • The SLR included up to 253 studies but the mapping study on PbD principles concerned 23 studies.
@@ 4.1 Research Questions (RQs) Definition
RQ1. How well does the literature on privacy in mobile apps cover the GDPR privacy concepts? This RQ explores the research landscape on mobile apps to identify the main topics in terms of GDPR privacy concerns. Drawing on the reference model (presented in Section 2), we answer RQ1 by mapping the topics investigated in existing work on mobile apps to the concepts in the conceptual model.
RQ2. What are the main objectives pursued when investigating GDPR privacy concerns? RQ2 explores the main objectives of primary studies that investigated GDPR privacy in mobile apps. The intuition behind this RQ is to understand the GDPR-related research challenges, which were acknowledged or addressed by existing work. As a result, RQ2 identifies remaining challenges that are currently under-explored and which can lead to potentially open directions for future research.
RQ3. What are the different types of research contribution? RQ3 explores the different types of research contribution made by the primary studies, distinguishing between technical and non-technical contributions, and analyzing the app stores and operating systems targeted in the literature. It provides a comprehensive view on the research contributions and further describes, in the case of technical solutions, the enabling technologies and the way such contributions were validated.
RQ4. Are the research artifacts accompanying studies on GDPR privacy concerns publicly available? This RQ sheds light on the publicly available artifacts that can be potentially leveraged in future research. RQ4 further reports on the licenses under which these artifacts have been shared.
@@ 4.2 Phase I: Paper Extraction
Ensuring the retrieval of a relatively representative sample of the existing literature is essential for conducting an SLR. This phase involves three steps: (1) selecting the databases or search engines to query, (2) refining the query keywords, and (3) retrieving the primary studies. Next, we elaborate on each step.
@@ 4.2.1 Databases Selection
For the search of papers in the different databases, we adopted common practices for and recommendations from the SLR literature [44, 43, 45]. Specifically, we retrieved primary studies from the following databases: ACM Digital Library [^2] https://dl.acm.org/, IEEE Xplore [^3] https://ieeexplore.ieee.org/, ScienceDirect [^4] https://www.sciencedirect.com/, Scopus [^5] https://www.scopus.com/, and SpringerLink [^6] https://link.springer.com/. These sources are typically selected in SLRs and systematic mapping studies [46, 47]. Additionally, we searched for papers in the following search engines or indexing systems: DBLP [^7] https://dblp.org/, Google Scholar [^8] https://scholar.google.com/, and Web of Science [^9] https://www.webofscience.com/wos/. The rationale behind including these additional sources is mainly to broaden the search results and ensure a complete coverage of the SE literature. Such indexing systems typically contain preprints or in-press papers that are not yet included in the regular databases mentioned above, and would be otherwise missed. We scoped our search against these indexing systems by restricting it to the period 2022–2023, as we assume the papers published before 2022 are already covered by the comprehensive set of results retrieved from the main databases. We acknowledge that restricting the scope to this two-year period may introduce a threat to validity, as it can lead to missing potentially relevant studies that not indexed in the primary databases. We elaborate on this validity consideration in Section 6.
4.2.2 Search Query Definition
@@
Initially, we tried different combinations of relevant keywords like “data protection”, “privacy”, and “mobile applications”. We observed that including some keywords would narrow the search results while others broaden them beyond the scope of our study, yielding largely irrelevant papers. We tested several queries across the different databases. As a result, we deemed the following keywords to be the best fit for this study: (“mobile app*” AND “data protection” AND “data privacy”). We queried all databases using the same query except for ScienceDirect, which does not allow using special characters such as wildcards(*). In this case, we used the following combination: (“mobile app” OR “mobile apps” OR “mobile application” OR “mobile applications”).
Finally, our final query is as follows:
‘‘mobile app’’ AND ‘‘privacy’’ AND ‘‘gdpr’’
@@ 4.2.3 Papers Retrieval
Table 2: Number of Papers Initially Retrieved per Database using the query ‘‘mobile apps’’ AND ‘‘privacy’’ AND ‘‘gdpr’’ and Remaining Papers after Filtering and Relevance Assessment.
Database Retrieval Results Filtering Results Relevant Papers
ACM Digital Library 181 61 23
IEEE Xplore 70 18 13
ScienceDirect 129 43 11
Scopus 26 5 5
SpringerLink 78 41 8
DBLP 15 0 0
Google Scholar 100 0 0
Web of Science 17 0 0
Summary 616 168 60
We applied the above criteria as filters across the selected databases. However, slight adaptations were necessary to accommodate the available options in each platform. For instance, criteria I1–I3 were fully applicable in ScienceDirect, Scopus, and SpringerLink. IEEE Xplore and the ACM Digital Library do not allow for explicit domain specification. This limitation is nonetheless acceptable since both platforms are focused on computer science and related scientific fields, and hence any irrelevant papers were subsequently excluded during the assessment step. In ACM Digital Library, journal (articles) and conference proceedings must be selected separately, requiring manual merging of the resulting lists. In DBLP, filters must be split by paper type and publication year, with results to be aggregated afterwards. Web of Science lacks options for specifying the domain, publication period, or paper type. Similarly, GoogleScholar does not support filtering by paper type or research domain.
The number of retrieved papers from each source are listed in the second column of Table 2. We note that the results presented in this study reflect the state of the art as of October 2023. Querying the regular databases resulted in a collection of 484 papers, distributed across the different databases as follows: 181 from ACM Digital Library, 70 from IEEE Xplore, 129 from ScienceDirect, 26 from Scopus, and 78 from SpringerLink. As noted earlier, we also queried indexing systems for additional potentially missing papers. Consequently, we retrieved 132 additional papers, including 100 papers from Google Scholar, 17 papers from the Web of Science, and 15 papers from DBLP. As for Google Scholar, we considered only the first 10 pages of results; we checked the results in the following pages and confirmed that they were not relevant to the objectives of our study. Following best practices, we did snowballing and further used the search string in dedicated venues. This process has not resulted in identifying any new papers that were not previously identified in our paper retrieval step. Our paper extraction phase resulted in a total of 616 papers, which were then passed on to the next phase for further analysis.
@@ 4.3.1 Paper filtering
The first step of the paper filtering consisted in removing the duplicates from the entire collection of 616 papers. To identify duplicates, we defined a case-insensitive Excel macro comparing the titles of the papers and determining possible duplicates. The output of this macro indicated which paper metadata details (particularly the title) were actually duplicates. This output was then manually vetted by the first author of this paper. As a result, we filtered out 204 papers which were deemed genuine duplicates, thereby reducing our collection to 412 papers.
@@
    Paper length: We defined a minimum paper size of six pages (including references). The rationale behind this threshold was to discard short papers such as position, vision, demo, new-idea, early-research-achievements papers with only preliminary or partial results, as commonly done in relevant literature [46, 47].
@@
    Publication venues: Following the guidelines of Kitchenham and Brereton [26], we filtered out papers that were not relevant for our study. Specifically, we excluded papers published in journals or conferences clearly unrelated to SE and privacy. In this way, we removed papers focusing on the social aspects of privacy/data protection or investigating specialized topics. In case of uncertainty (e.g., when a paper title appeared relevant), we retained the venue and related paper(s) for further analysis (see below). Examples of excluded venues are the International Journal of Medical Informatics, Blockchain: Research and Applications, and the EAI International Conference on Smart Objects and Technologies for Social Good.
At the end of the paper filtering, we were left with 168 papers; the details of their sources are provided in the third column of Table 2. The venues (36 journals and 24 conferences) and the respective number of papers (105 journal papers and 63 conference papers) are summarized in column N of Tables 3 and 4. These papers were then processed through the relevance assessment step.
4.3.2 Relevance Assessment:
We analyzed the remaining 168 papers and identified those relevant for our SLR. The first three authors collectively participated in this process. Driven by the motivation of our SLR, the relevance of the papers was assessed according to their main focus being privacy.
Therefore, we considered a paper relevant if:
  • I4
    The paper clearly discussed GDPR-relevant privacy concerns and how these were handled in mobile apps. For instance, a paper that discusses users practices in mobile apps, such as subscription cancellation and asking for refunds, or users providing inaccurate information were considered as not relevant to our SLR. Similarly, papers discussing dark patterns that attempt to manipulate users’ behaviors are not relevant to our SLR. While these papers could be relevant to privacy in a broader sense, they are less informative about how the handling and processing of personal data were addressed in mobile apps.
The relevance assessment step spanned three iterations, described below.
@@ 4.3.2 Relevance Assessment:
The output of this step is the final set of 60 papers (30 journal papers and 30 conference papers), as indicated in column N of Tables 3 and 4), corresponding to 35.7% of the 168 analyzed papers, on which we conducted our data extraction phase.
Table 3: Overview of journal articles, assessed (column N) and deemed relevant (column N) in this SLR.
[t] ID Journal Name N N J1 Journal of Computer Science and Technology 1 0 J2 Journal of Engineering and Applied Science 1 0 J3 Mobile Networks and Applications 1 0 J4 Internet Interventions 1 1 J5 Annals of Telecommunications 1 0 J6 Information and Software Technology 3 1 J7 Wireless Networks 1 1 J8 Computers & Security 12 4 J9 Soft Computing 1 0 J10 Journal of Systems and Software 3 1 J11 Complex & Intelligent Systems 1 0 J12 Frontiers of Computer Science 2 1 J13 Journal of Information Security and Applications 3 0 J14 International Journal of Information Technology 1 0 J15 Human-Centric Intelligent Systems 1 0 J16 Computer Law & Security Review 16 4 J17 AI and Ethics 2 0 J18 Software and Systems Modeling 1 1 J19 Intelligent Systems with Applications 1 0 J20 Computing 2 1 J21 Information Processing & Management 1 0 J22 Empirical Software Engineering 5 2 J23 Computer Fraud & Security 1 0 J24 Artificial Intelligence and Law 1 0 J25 Information Systems 1 0 J26 International Cybersecurity Law Review 1 0 J27 AI & Society 8 1 J28 International Journal of Information Security 4 1 J29 Data Science and Management 1 0 J30 Ethics and Information Technology 3 0 J31 Frontiers of Information Technology & Electronic Engineering 1 0 J32 ACM Transactions on Computer-Human Interaction 10 7 J33 Personal and Ubiquitous Computing 3 1 J34 IEEE Transactions on Software Engineering 0 2 J35 ACM Transactions on Software Engineering Methodology 2 1 J36 Proceedings of the ACM on Computer-Human Interaction 8 0 Total 105 30
Table 4: Overview of conference papers, assessed (column N) and deemed relevant (column N) in this SLR.
[t] ID Venue N N V1 International Conference on Mining Software Repositories 1 0 V2 IEEE International Conference on Software Analysis, Evolution and Reengineering 1 1 V3 International Conference on Availability, Reliability and Security 3 1 V4 Annual ACM Symposium on User Interface Software and Technology 1 0 V5 International Conference on Computer Systems and Technologies 1 0 V6 Annual Computer Security Applications Conference 1 1 V7 The Web Conference 1 1 V8 ACM Web Conference 4 1 V9 Annual International Conference on Mobile Computing and Networking 1 0 V10 Annual IEEE/IFIP International Conference on Dependable Systems and Network 1 1 V11 USENIX Security Symposium 2 2 V12 IEEE Symposium on Security and Privacy 1 0 V13 Annual Computers, Software, and Applications Conference 1 0 V14 ACM SIGSAC Conference on Computer and Communications Security 4 2 V15 ACM Conference on Security and Privacy in Wireless and Mobile Networks 1 1 V16 IEEE/ACM International Conference on Automated Software Engineering 6 5 V17 CHI Conference on Human Factors in Computing Systems 10 1 V18 IEEE International Requirements Engineering Conference 3 3 V19 ACM SIGSOFT International Symposium on Software Testing and Analysis 1 0 V20 IEEE European Symposium on Security and Privacy Workshops 4 2 V21 ACM/IEEE International Conference on Software Engineering 6 6 V22 IEEE Annual Computer Software and Applications Conference 1 1 V23 IEEE International Conference on Trust, Security and Privacy in Computing and Communications 1 0 V24 International Conference on Mobile and Ubiquitous Multimedia 1 0 V25 Workshop on Privacy in the Electronic Society 4 1 Total 63 30